if o2 is bubbled through water at 393k, how many milimoles of o2 gas would be resolved in one litreof water? assume that o2 exerts of pressure 0.95 bar?
1
Expert's answer
2018-08-22T05:06:37-0400
Question:

if o2 is bubbled through water at 393k, how many milimoles of o2 gas would be resolved in one litreof water? assume that o2 exerts of pressure 0.95 bar?

T=393K
n(O2)-?
V=1 l = 0.001 m3
P = 0.95 bar

Solution:

PV = nRT , n=PV/RT
R = 8.31 J/mol*K
1 bar = 100 kPa, 0.95 bar = 95 kPa = 95000 Pa
Then, n = (95000*0.001)/(8.31*393) = 0.029 mol = 29 millimoles
Answer: 29 millimoles of O2